A helicopter takes off vertically. It starts from rest at time t=0, and has a constant
upwards acceleration of 0.40 m/s2. The mass of the helicopter is 14 000 kg. The
acceleration of free fall is 10 m/s2
.
1
Expert's answer
2020-06-09T13:19:29-0400

The speed of the helicopter at time 20 s will be


"v(t)=v_0+at,\\
v(20)=0+0.40\cdot20=8.0\text{ m/s}."

The resultant force acting on the helicopter:


"F=ma=14000\cdot0.40=5600\text{ N}."